# Psychiatry-Bench: A Multi-Task Benchmark for LLMs in Psychiatry

### v1

@@ 4. Provide a single numeric score between 0 and 100:
    3. Evaluate the response on:
    Definition Accuracy & Completeness: Are all core features of the concept present and
    correctly described?
    Clarity & Precision: Is the explanation clear, unambiguous, and clinically precise?
    Depth of Explanation: Does it include relevant examples or elaborations that demonstrate
    understanding?
@@
Concept to Define : {Answer}
@@
Template 11: Sequential Question Answering Prediction Prompt
Patient History: {History}
Task: Based on the provided history, answer the following five clinical questions. Use the
same format and level of detail shown in the example. Make sure to include preferred and
alternative diagnoses where applicable, and support your answers with reasoning drawn fromthe
case.
@@
Q2: What information in the history supports the diagnosis, and what other information would
    help to confirm it?
@@
    (Cover location-based options, physical treatments, psychological support, and social
    considerations. Mention use of Mental Health Act if relevant.)
@@
    (Discuss short-term and long-term outlook. Include good and poor prognostic indicators
    based on patient specifics.)
@@
Q6: based on the provided History ,which class does this diagnosis belong to according to
    DSM-5 ?
Template 12: Sequential Question Answering Evaluation Prompt
@@
SYSTEM: You are a mental health case response evaluator. Your task is to assess six clinical
@@
responses from a trainee clinician. Each response corresponds to a specific psychiatric
question.
@@
For each question, compare the candidate’s answer to the expert reference answer and provide a
numeric score based on how well the candidate captures the clinical reasoning and content.
    • Assign a similarity score between 0 and 100:
        - 0: Text A does not contain the correct diagnoses or significantly misrepresents key
        clinical reasoning compared to Text B.
        - 50: Text A identifies some key diagnoses but either misses important differentials or
        partially misrepresents clinical reasoning compared to Text B.
        - 100 means Text A correctly captures all essential diagnoses, differential diagnoses,
        and clinical reasoning as stated in Text B.
    • Use the following criteria to guide your scoring:
        - Presence of the correct main diagnosis (40 points).
        - Inclusion and accuracy of differential diagnoses (30 points).
    • Provide ONLY the numeric similarity score (0-100) without any additional explanation or
    justification.
@@
Template 13: Multi-stage prompting strategies for objective assessment items. (a) MCQ Prompt 1 directs the model to choose one answer without explanation; (b) MCQ Prompt 2 adds a brief analytical step over patient history; (c) EMI Prompt 1 handles extended matching items where options may be reused or omitted; (d) EMI Prompt 2 blendsn-based EMIs with standard MCQs, requiring only the answer letters.
@@
Answer the question using one of the given choices.
Do not provide any explanation for your choice. Indicate your choice with a single English
@@
letter. (Provide only one answer; Multiple answers could be provided.)
History:{History}
@@
Analyze the multiple-choice question (MCQ) and select the best answer.
Assess each option logically based on relevant psychological history.
Do not provide any explanation for your choice. Indicate your choice with a single English
letter. (Provide only one answer; Multiple answers could be provided.)
History:{History}
Question:{Question}
Choices:{Choices}
MCQ Prompt 2
Analyze the extended matching items (EMI) and select the best answer. each option (denoted a,
b, c, etc.)
might be used once, more than once, or not at all. Assess each option logically. Please don’t
give any explanation for your choice.
@@
Analyze the multiple-choice question (MCQ) and select the best answer.
Assess each option logically.
Do not provide any explanation for your choice. Indicate your choice with a single English
letter. (Provide only one answer; Multiple answers could be provided.)
EMI Prompt 2
Template 14: Point-based similarity EMI prompting templates: (a) a straightforward "header-required" template that simply presents the fields, (b) an analytical variant that adds an instruction to assess each option logically before choosing.
Answer the question using one of the given choices.
Do not provide any explanation for your choice. Indicate your choice with a single English
letter. (Provide only one answer; Multiple answers could be provided.)
@@
Analyze the multiple-choice question (MCQ) and select the best answer.
Assess each option logically .
Do not provide any explanation for your choice. Indicate your choice with a single English
letter. (Provide only one answer; Multiple answers could be provided.)
Header:{Header}
Required:{Required}
Question:{Question}
Choices:{Choices}
(b) EMI Separated Prompt 2
Template 15: Prompt generated by GPT-4.5 for point-based similarity scoring of clinical reasoning responses.
I will provide two texts: Text A (LLM’s response) and Text B (True answer). Evaluate their
similarity based strictly on the presence and accuracy of key diagnoses, differential
diagnoses, and clinical reasoning. Follow these guidelines:
    • Assign a similarity score between 0 and 100:
        - 0: Text A does not contain the correct diagnoses or significantly misrepresents key
        clinical reasoning compared to Text B.
        - 50: Text A identifies some key diagnoses but either misses important differentials or
        partially misrepresents clinical reasoning compared to Text B.
        - 100 means Text A correctly captures all essential diagnoses, differential diagnoses,
        and clinical reasoning as stated in Text B.
    • Use the following criteria to guide your scoring:
        - Presence of the correct main diagnosis (40 points).
        - Inclusion and accuracy of differential diagnoses (30 points).
    • Provide ONLY the numeric similarity score (0-100) without any additional explanation
    or justification.
Template 16: Prompt generated by Gemini for Llama 3 70B evaluation.
SYSTEM: You are a medical diagnosis evaluator. Your task is to compare two medical text
summaries and assess the degree to which the first text (the "Candidate Answer") captures the
essential diagnostic information contained in the second text (the "Reference Answer"). Focus
on the core medical meaning and the completeness of the diagnoses, not on the writing style or
specific wording.
USER: INSTRUCTIONS:
1. Read the "Reference Answer" carefully. Identify the key diagnoses (both primary and
differential) and any crucial supporting information (e.g., symptoms, test results, ruled-out
disorderss).
2. Read the "Candidate Answer". The "Candidate Answer" will be a concise list of diagnoses and
key findings.
    • Diagnostic Accuracy: Does the "Candidate Answer" correctly identify the most likely
    diagnosis (if one is clearly indicated in the "Reference Answer")?
    • Diagnostic Completeness: Does the "Candidate Answer" include all of the major
    differential diagnoses mentioned in the "Reference Answer"? It is acceptable if the
    "Candidate Answer" uses slightly different terminology as long as the underlying medical
    concept is the same.
    • Supporting Information: Does the "Candidate Answer" include the key symptoms, test
    results, or other evidence that support the diagnoses? It doesn’t need to include every
    detail, but it should capture the most important ones. It’s ok if this supporting
    information is implicitly present in the reference answer.
    • Exclusion of Irrelevant Information: Do not penalize the "Candidate Answer" for including
    additional information that is not present in the "Reference Answer", as long as that
    information is medically accurate and relevant to the case.
    • groundance does not affect the score: Do not penalize a response for being redundant.
    • 0: The "Candidate Answer" has no meaningful overlap with the "Reference Answer" in terms
    of diagnosis or supporting information. It’s completely wrong.
    • 50: The "Candidate Answer" identifies some relevant diagnoses or supporting information,
    but it misses major elements or includes significant inaccuracies.
    • 75: The "Candidate Answer" correctly identifies the primary diagnosis (if applicable) and
    most of the important differential diagnoses and supporting information. It may miss some
    minor details or include some slight inaccuracies.
    • 90: The "Candidate Answer" is very close to the "Reference Answer" in terms of core
    diagnostic meaning and completeness. It may have minor omissions or differences in
    emphasis, but it captures almost all of the essential information.
    • 100: The "Candidate Answer" is a perfect match in terms of core diagnostic meaning. It
    includes all of the key diagnoses and supporting information from the "Reference Answer",
    even if the wording or organization is different.
Here is an example for better understanding:
Reference Answer:
Differential diagnosis: 1. Delirium 2. Substance intoxication 3.Substance withdrawal delirium
4. Mental retardation
Candidate Answer:
- Primary Diagnosis: Alzheimer’s_Disease
- Differential Diagnoses: Vascular_Dementia, Lewy_Body_Dementia, Frontotemporal_Dementia
ASSISTANT: 75
Here is another example:
Reference Answer:
Alcohol Use Disorder (Severe)
Candidate Answer:
- Primary Diagnosis: Alcohol_Use_Disorder (Mild)
- Key Findings: History of heavy alcohol use, withdrawal_symptoms.
ASSISTANT: 50
Now, evaluate the following:
Template 17: Treatment Prediction Prompt
You are a psychiatrist. Based on the patient’s clinical history, answer the following question
with a medically appropriate and individualized treatment recommendation.
Guidelines
- Base your treatment recommendation primarily on the Patient’s History. Use information from
the Mental and Physical Examinations only if it directly supports your treatment choice (if
found).
- Your response should be clinically sound, concise, and focused on the patient’s specific
presentation.
- Justify your recommendation using evidence from the case. Do not include general advice or
unrelated commentary.
- Respond in full, professional sentences.
Patient’s History :{History}
Mental State Examination : {Mental State Examination}
Physical Examination : {Physical Examination}
Question : {Question}
Template 18: Treatment Evaluation Prompt
SYSTEM: You are a medical treatment evaluator. Your task is to compare two medical treatment
summaries and assess the degree to which the first text (the "Candidate Answer") captures the
essential therapeutic interventions described in the second text (the "Reference Answer").
Focus on the medical appropriateness and completeness of the treatment plan, *not on the
writing style or specific wording.
USER:
INSTRUCTIONS:
1. Read the "Reference Answer" carefully. Identify the key components of the treatment plan,
including primary therapeutic interventions, medications, behavioral strategies, or other
relevant clinical actions.
2. Read the "Candidate Answer". The "Candidate Answer" will be a concise summary of proposed
treatments and strategies.
     Treatment Accuracy: Does the "Candidate Answer" correctly identify the key treatment or
    intervention as outlined in the "Reference Answer"?
     Treatment Completeness: Does the "Candidate Answer" include *all of the major treatment
    components mentioned in the "Reference Answer"? Slightly different terminology is
    acceptable as long as the clinical *concept is preserved.
     Supporting Rationale: Does the "Candidate Answer" capture the reasoning or context for
    the treatment (e.g., why a specific method or medication is chosen)? It doesn’t need to
    list every detail, but should reflect the main justification.
     Exclusion of Irrelevant Information: Do *not penalize the "Candidate Answer" for
    including additional, medically accurate and relevant information not found in the
    "Reference Answer".
     groundance does not affect the score: Do not penalize a response for being redundant.
     0: The "Candidate Answer" has no meaningful overlap with the "Reference Answer" in terms
    of treatment or rationale. It is completely inappropriate or unrelated.
     50: The "Candidate Answer" identifies *some relevant treatment components or
    strategies, but misses key parts or includes significant inaccuracies.
     75: The "Candidate Answer" correctly identifies the primary treatment and most important
    additional strategies or reasoning. Minor omissions or slight inaccuracies are acceptable.
     90: The "Candidate Answer" is very close to the "Reference Answer" in terms of core
    treatment strategies and rationale. It may have minor differences in emphasis or detail.
     100: The "Candidate Answer" is a perfect match in terms of treatment content and
    rationale. It includes all key interventions and supporting reasoning from the "Reference
    Answer", even if the wording or structure is different.
Here is an example for better understanding:
Reference Answer:
Treatment includes structured desensitization therapy and cognitive behavioral therapy (CBT).
Desensitization is done through exposure to a graded hierarchy of fear situations, supported
by relaxation techniques and possibly hypnotherapy. Treatment of the current abscess under
sedation is required first, followed by referral to a specialist dentist and CBT led by a
mental health professional.
Candidate Answer:
- Suggested Treatment: CBT for dental phobia
- Steps: Desensitization with imagery and real-life exposure, use of sedation for abscess
treatment, follow-up with a specialist dentist
ASSISTANT: 100
Here is another example:
Reference Answer:
CBT and desensitization therapy are the main treatments. Use of sedation is recommended
initially to manage acute dental issues. Follow-up should include exposure-based therapy and
psychoeducation.
Candidate Answer:
- Suggested Treatment: Use of medication for phobia and pain
- Notes: Patient should be given anti-anxiety medications and then referred to a dentist
ASSISTANT: 50
Now, evaluate the following:
Reference Answer:
{Answer}
Candidate Answer:
{response}
@@ 10.5 Classification Mapping
Template 19: Specific Disorders List
@@
"Language Disorder", "Speech Sound Disorder", "Childhood-Onset Fluency Disorder (Stuttering)",
@@
"Unspecified neurodevelopmental disorders","Delusional Disorder", "Brief Psychotic Disorder",
@@
"Other Specified Bipolar and Related Disorder", "Unspecified Bipolar and Related Disorder",
@@
"Separation Anxiety Disorder", "Selective Mutism", "Specific Phobia","Social Anxiety Disorder",
@@
"Unspecified Anxiety Disorder", "Obsessive-Compulsive Disorder", "Body Dysmorphic Disorder",
@@
"Psychological Factors Affecting Other Medical disorderss", "Factitious Disorder",
@@
Template 20: Disorder Categories List
["Intellectual Developmental Disorders","Communication Disorders",
@@
"Specific Learning Disorder","Motor Disorders","Tic Disorders",
"Schizophrenia Spectrum and Other Psychotic Disorders",
"Bipolar and Related Disorders","Depressive Disorders","Anxiety Disorders",
"Obsessive-Compulsive and Related Disorders","Trauma- and Stressor-Related Disorders",
"Dissociative Disorders","Somatic Symptom and Related Disorders",
"Feeding and Eating Disorders","Elimination Disorders","Sleep-Wake Disorders",
@@
"Breathing-Related Sleep Disorders","Parasomnias","Sexual Dysfunctions","Gender Dysphoria",
"Disruptive, Impulse-Control, and Conduct Disorders",
"Substance-Related and Addictive Disorders","Neurocognitive Disorders",
"Personality Disorders","Medication-induced Movement Disorders","Paraphilic Disorders"]

### v2

@@ 1. Read the "Answer" this is the clinical concept or term to define
    3. Evaluate the response on:Definition Accuracy & Completeness: Are all core features of the concept present
    and correctly described? Clarity & Precision:Is the explanation clear, unambiguous, and clinically precise?
    Depth of Explanation: Does it include relevant examples or elaborations that demonstrate understanding?
@@ 4. Provide a single numeric score between 0 and 100:
Reference Answer : {Answer}
@@
Template 13: Sequential Question Answering Prediction Prompt
The Patient’s History: {History}
Task: Based on the provided history, answer the following Six clinical questions. Use the same format and
level of detail shown in the example. Make sure to include preferred and alternative diagnoses where applicable
,and support your answers with reasoning drawn fromthe case.
@@
Q2: What information in the history supports the diagnosis, and what other information
would help to confirm it?
@@
    (Cover location-based options, physical treatments, psychological support,
    and social considerations. Mention use of Mental Health Act if relevant.)
@@
    (Discuss short-term and long-term outlook. Include good and poor prognostic
    indicators based on patient specifics.)
@@
Q6: based on the provided History ,which class does this diagnosis belong to according
to DSM-5-TR ?
Template 14: Sequential Question Answering Evaluation Prompt (Part 1)
@@
SYSTEM: You are a mental health case response evaluator. Your task is to assess six clinical responses
@@
from a trainee clinician. Each response corresponds to a specific psychiatric question. For each question,
@@
compare the candidate’s answer to the expert reference answer and provide a numeric score based on how well
the candidate captures the clinical reasoning and content.
Q1: What is the likely differential diagnosis?
- Task Type: Diagnosis
- Focus: Diagnostic reasoning and clinical hypotheses.
- Scoring System:
    0: Completely off-topic, unrelated, or clinically incorrect.
    50: Some valid elements of the diagnosis are present, but many key elements are missing or poorly reasoned.
    100: Fully matches the core diagnosis and rationale.
Concept to Define: {Answer for Question 1}
Candidate Response: {Response for Question 1}
Q2: What information in the history supports the diagnosis, and what other information would help to confirm
it?
- Task Type: History
- Focus: Presenting symptoms, clinical history, and psychosocial context.
- Scoring System:
    0: Completely off-topic, unrelated, or clinically incorrect.
    50: Some valid elements of the history are present, but many key elements are missing or poorly reasoned.
    100: Fully matches the core history and rationale.
Concept to Define: {Answer for Question 2}
Candidate Response: {Response for Question 2}
Q3: What might the important aetiological factors be?
- Task Type: Aetiology
- Focus: Biological, psychological, and social contributing factors.
- Scoring System:
    0: Completely off-topic, unrelated, or clinically incorrect.
    50: Some valid elements of the aetiology are present, but many key elements are missing or poorly reasoned.
    100: Fully matches the core aetiology and rationale.
Concept to Define: {Answer for Question 3}
Candidate Response: {Response for Question 3}
Q4: What treatment options are available?
- Task Type: Treatment
- Focus: Clinical interventions including medication, therapy, and lifestyle recommendations.
- Scoring System:
    0: Completely off-topic, unrelated, or clinically incorrect.
    50: Some valid elements of the treatment are present, but many key elements are missing or poorly reasoned.
    100: Fully matches the core treatment and rationale.
Concept to Define: {Answer for Question 4}
Candidate Response: {Response for Question 4}
Template 15: Sequential Question Answering Evaluation Prompt (Part 2)
Q5: What is the prognosis in this case?
- Task Type: Prognosis
- Focus: Outcome prediction, expected recovery, and risk factors.
- Scoring System:
    0: Completely off-topic, unrelated, or clinically incorrect.
    50: Some valid elements of the prognosis are present, but many key elements are missing or poorly reasoned.
    100: Fully matches the core prognosis and rationale.
Concept to Define: {Answer for Question 5}
Candidate Response: {Response for Question 5}
Q6: Based on the provided history, which class does this diagnosis belong to according to DSM-5 or ICD-10?
- Task Type: Classification
- Focus: Diagnostic nosology and classification system alignment.
- Scoring System:
    0: Completely off-topic, unrelated, or clinically incorrect.
    50: Some valid elements of the classification are present, but many key elements are missing or poorly
    reasoned.
    100: Fully matches the core classification and rationale.
Concept to Define: {Answer for Question 6}
Candidate Response: {Response for Question 6}
Now:
- Return your scores only in this JSON format, with numeric values only.
- Then, sum the six scores and divide by 6 to compute the final score.
- Return only the final score as a single number (no labels, no JSON, no extra text).
  "Q1": ___,
  "Q2": ___,
  "Q3": ___,
  "Q4": ___,
  "Q5": ___,
  "Q6": ___
Final Score: ___
@@
Template 16: Prompting strategies for objective assessment items. (a) MCQ Prompt 1 directs the model to choose one answer without explanation; (b) EMI Prompt 2 blends scenario-based extended matching items (EMIs) with standard MCQs, requiring only the answer letters.
@@
Answer the question using one of the given choices.Do not provide any explanation for your choice.
@@
Indicate your choice with a single English letter.Provide only one answer; Multiple answers could be provided
The Patient’s History:{History}
@@
Analyze the multiple-choice question (MCQ) and select the best answer.Assess each option logically based on
relevant psychological history.Do not provide any explanation for your choice. Indicate your choice with
a single English letter. (Provide only one answer; Multiple answers could be provided.)
The Patient’s History:{History}
@@
Template 17: Point-based similarity EMI prompting templates: (a) a straightforward "header-required" template that simply presents the fields.
Answer the question using one of the given choices.Do not provide any explanation for your choice. Indicate
your choice with a single English letter. (Provide only one answer; Multiple answers could be provided.)
@@ 1.5 Classification Mapping
This subsection details the standardized diagnostic taxonomy employed for the classification task, which was explicitly derived from the DSM-5-TR to ensure terminological precision. To facilitate rigorous multi-label evaluation, two distinct reference lists were curated and validated by a licensed psychiatrist : a granular list of Specific Disorders  18, encompassing detailed diagnoses such as "Bipolar I Disorder," and a high-level list of Disorder Categories  19, which groups conditions into broader classes like "Depressive Disorders." These mappings serve as the ground truth for evaluation, requiring models to align their outputs with authoritative psychiatric nosology rather than relying on ambiguous or unstructured free-text descriptions
Template 18: Specific Disorders List
@@
"Language Disorder", "Speech Sound Disorder",
"Childhood-Onset Fluency Disorder (Stuttering)",
@@
"Unspecified neurodevelopmental disorders","Delusional Disorder",
"Brief Psychotic Disorder",
@@
"Other Specified Bipolar and Related Disorder",
"Unspecified Bipolar and Related Disorder",
@@
"Separation Anxiety Disorder", "Selective Mutism", "Specific Phobia",
"Social Anxiety Disorder",
@@
"Unspecified Anxiety Disorder", "Obsessive-Compulsive Disorder",
"Body Dysmorphic Disorder",
@@
"Psychological Factors Affecting Other Medical disorders", "Factitious Disorder",
@@
Template 19: Disorder Categories List
["Intellectual Developmental Disorders", "Communication Disorders", "Autism Spectrum Disorder",
@@
"Attention-Deficit/Hyperactivity Disorder", "Specific Learning Disorder", "Motor Disorders",
"Tic Disorders","Schizophrenia Spectrum and Other Psychotic Disorders", "Bipolar and Related Disorders",
"Depressive Disorders", "Anxiety Disorders","Obsessive-Compulsive and Related Disorders",
"Trauma- and Stressor-Related Disorders", "Dissociative Disorders", "Somatic Symptom and Related Disorders",
"Feeding and Eating Disorders", "Elimination Disorders", "Sleep-Wake Disorders", "Parasomnias",
@@
"Sexual Dysfunctions","Breathing-Related Sleep Disorders", "Gender Dysphoria", "Disruptive,
Impulse-Control, and Conduct Disorders", "Substance-Related and Addictive Disorders",
"Neurocognitive Disorders", "Personality Disorders", "Medication-Induced Movement Disorders",
"Paraphilic Disorders"]
Template 20: Prompt generated by GPT-4.5 for point-based similarity scoring of clinical reasoning responses.
I will provide two texts: Text A (LLM’s response) and Text B (True answer). Evaluate their similarity based
strictly on the presence and accuracy of key diagnoses, differential diagnoses, and clinical reasoning.
Follow these guidelines:
    • Assign a similarity score between 0 and 100:
        - 0: Text A does not contain the correct diagnoses or significantly misrepresents key clinical reasoning
        compared to Text B.
        - 50: Text A identifies some key diagnoses but either misses important differentials or partially
        misrepresents clinical reasoning compared to Text B.
        - 100 means Text A correctly captures all essential diagnoses,
        differential diagnoses,and clinical reasoning as stated in Text B.
    • Use the following criteria to guide your scoring:
        - Presence of the correct main diagnosis (40 points).
        - Inclusion and accuracy of differential diagnoses (30 points).
    • Provide ONLY the numeric similarity score (0-100) without any additional explanation or justification.
Template 21: Prompt generated by Gemini for Llama 3 70B evaluation.
SYSTEM: You are a medical diagnosis evaluator. Your task is to compare two medical text summaries and assess
the degree to which the first text (the "Candidate Answer") captures the essential diagnostic information
contained in the second text (the "Reference Answer"). Focus on the core medical meaning and the completeness
of the diagnoses, not on the writing style or specific wording.
USER:
INSTRUCTIONS:
1. Read the "Reference Answer" carefully. Identify the key diagnoses (both primary and differential) and any
crucial supporting information (e.g., symptoms, test results, ruled-out disorders).
2. Read the "Candidate Answer". The "Candidate Answer" will be a concise list of diagnoses and key findings.
    • Diagnostic Accuracy: Does the "Candidate Answer" correctly identify the most likely diagnosis
    (if one is clearly indicated in the "Reference Answer")?
    • Diagnostic Completeness: Does the "Candidate Answer" include all of the major differential diagnoses
    mentioned in the "Reference Answer"? It is acceptable if the "Candidate Answer" uses slightly different
    terminology as long as the underlying medical concept is the same.
    or other evidence that support the diagnoses? It doesn’t need to include every detail, but it should capture
    the most important ones.It’s ok if this supporting information is implicitly present in the reference answer
    • Exclusion of Irrelevant Information:
    Do not penalize the "Candidate Answer" for including additional
    information that is not present in the "Reference Answer", as long as that information is medically accurate
    and relevant to the case.
    • groundance does not affect the score:Do not penalize a response for being redundant
    • 0: The "Candidate Answer" has no meaningful overlap with the "Reference Answer" in terms of diagnosis or
    supporting information. It’s completely wrong.
    • 50: The "Candidate Answer" identifies some relevant diagnoses or supporting information, but it misses
    major elements or includes significant inaccuracies.
    • 75: The "Candidate Answer" correctly identifies the primary diagnosis
    (if applicable) and most of the important differential diagnoses and supporting information. It may miss
    some minor details or include some slight inaccuracies.
    • 90: The "Candidate Answer" is very close to the "Reference Answer" in terms of core diagnostic meaning and
    completeness. It may have minor omissions or differences in emphasis, but it captures almost all of
    the essential information.
    • 100: The "Candidate Answer" is a perfect match in terms of core diagnostic meaning. It includes all of
    the key diagnoses and supporting information from the "Reference Answer",even if the wording
    or organization is different.
Here is an example for better understanding:
Reference Answer:
Differential diagnosis: 1. Delirium 2. Substance intoxication 3.Substance withdrawal
delirium 4. Mental retardation
Candidate Answer:
- Primary Diagnosis: Alzheimer’s_Disease
- Differential Diagnoses: Vascular_Dementia,
Lewy_Body_Dementia, Frontotemporal_Dementia
ASSISTANT: 75
Here is another example:
Reference Answer:
Alcohol Use Disorder (Severe)
Candidate Answer:
- Primary Diagnosis: Alcohol_Use_Disorder (Mild)
- Key Findings: History of heavy alcohol use, withdrawal_symptoms.
ASSISTANT: 50
Now, evaluate the following:
Reference Answer : {Answer}
Candidate Response: {Response}